Find the slope of the line containing the points (7, 9) and (1, 1)
Answer by Boreal(15235)

slope = changes in y/changes in x
may go in either direction, but MUST be consistent
left to right 9-1/7-1
=8/6 or 4/3
Do it the other way
1-9/1-7 -8/-6 =4/3
Be careful when the second number is negative
1,-1) (3,4)
y=4 - (-1)/3-1 4-(-1)=5
5/2